• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 21:38:09 on December 11, 2008

    Entry number 253031

    This entry is a followup to: 253011

    keyword=Compton asymmetry really did flip after HWP inserted

    I'm once more measuring asymmetries in the Compton detector after ~8 hours without Compton spectra.

    Figure 1 shows the asymmetries around noon today, before the HWP was inserted. We see an asymmetry that's a bit over 3% at the Compton edge. At the Compton edge it's positive for left photon polarization and negative for right photon polarization.

    Figure 2 shows the asymmetries from a run that finished at 21:12 this evening, with same voltage on the PMT for the photon detector, same gain settings, same thresholds, etc. We now see an asymmetry a bit under 4% at the Compton edge, but now it's positive for right photon polarization and negative for left.

    This is exactly what we expect, since a half-wave plate was inserted this evening between the two runs. This would flip the sign of the asymmetry.

    I would not put a lot of weight on the precise asymmetry numbers here, as these spectra are not fantastic and the error bars are large. But I think we can say that the beam is still polarized, and that the insertion of the HWP had the expected effect.

    It is still a mystery why the Compton spectrum disappeared earlier today. We are pretty sure it was a detector or DAQ problem and that there was nothing wrong with the beam polarization.

    A copy of this log entry has been emailed to: jiang@jlab.org, nanda@jlab.org, gbfranklin@cmu.edu, bquinn@cmu.edu



    Figure 1



    Figure 2